OpenStudy (anonymous):

Solve. 3x^2 + 12x = 3

OpenStudy (anonymous):

3x^2+12x-3=0 x^2+4x-1=0 (x+2)^2-5=0 follow that

OpenStudy (anonymous):

\[x=(-b \pm \sqrt{b^2-4ac})/2a\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

welcome :)